OpenStudy (anonymous):

what is (1/2)mv^2 = mgh ?? where this come from?

OpenStudy (anonymous):

hi

OpenStudy (roadjester):

(1/2)mv^2 is the definition of kinetic energy mgh is the definition of gravitational potential energy

OpenStudy (anonymous):

Try this http://www.wolframalpha.com/input/?i=1%2F2%29mv%5E2+%3D+mgh

OpenStudy (anonymous):

Does that help?

OpenStudy (anonymous):

i was trying to solve a physics problem, but then someone posted a solution starting with this equation (1/2)mv^2 = mgh

OpenStudy (anonymous):

box with given mass is projected up and he used this.

OpenStudy (roadjester):

can you type the question?

OpenStudy (anonymous):

A block of mass M is projected up a frictionless inclined plane with a speed VO. The angle of incline is given as theta.

OpenStudy (anonymous):

How am I supposed to know which equation to use..?

OpenStudy (anonymous):

like someone used (1/2)mv^2 = mgh?

OpenStudy (roadjester):

when you say projected, you you mean it was fired up an incline?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok so basically reason why I was confused with that equation is theres no such equation (1/2)mv^2 = mgh in my text book.

OpenStudy (roadjester):

that's because there wouldn't be.

OpenStudy (roadjester):

this equation is equating kinetic energy with potential energy

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

okok I will work on it!

OpenStudy (roadjester):

what you have to understand though is that if there is an external force acting on the system, then you can ADD potential and kinetic energy;

OpenStudy (roadjester):

if you are working within an isolated system, then energy must be conserved. have you learned about conservation of energy?

OpenStudy (anonymous):

Yes I do

OpenStudy (roadjester):

hi @douglaswinslowcooper could you help me a bit? I'm not sure how much to talk about

OpenStudy (anonymous):

This meant that the kinetic energy gained (1/2)m v^2 came from the potential energy lost m g h.

OpenStudy (anonymous):

This equation was from the Newton's Law subject

OpenStudy (roadjester):

ok and building on what @douglaswinslowcooper just said you lose potential energy, but it is converted into kinetic energy you do not however see this as an equation in the book because it is expected that you learn that energy is conserved

OpenStudy (anonymous):

oh i see

OpenStudy (anonymous):

I'm cleared now

OpenStudy (anonymous):

For an object falling due to gravity, the sum of potential and kinetic energy remain constant (1/2)mv^2 + m g h. If you define g as <0 or>0 you will have a + or - problem to pay attention to

OpenStudy (anonymous):

Ok I will work on it thank you so much